### v4

# Coloring Grids Avoiding Bicolored Paths

Abstract.
The star chromatic number on a graph is the minimum number of colors in a proper vertex coloring forbidding any $P_{4}$ with two colors (bicolored). This problem was introduced by Grünbaum (1973) together with the acyclic coloring of graphs, where bicolored cycles are avoided. In this paper, we study a generalization of this problem, by considering proper vertex coloring on graphs forbidding bicolored paths of a fixed length, which was initially discussed by Alon, McDiarmid, and Reed (1991). Here, we study this problem on products of two paths. We show that at least 4 colors are needed to properly color the product of paths, $P_{m}\square P_{n}$, avoiding a bicolored $P_{k},$ unless n<k2n<k-2 or m<k2.m<k-2. With this result, the above question is settled for all $k$ on 2-dimensional grids.
Key words and phrases:
star coloring, acyclic coloring, bicolored path, grid, graph product
1. Introduction
For a graph $G$, $\chi(G)$ denotes the chromatic number of $G$, that is the minimum number of colors needed to color $V(G)$ without a monochromatic edge. The star coloring problem on a graph $G$ asks to find the minimum number of colors in a proper coloring forbidding a bicolored (2-colored) $P_{4},$ called the star-chromatic number $\chi_{s}(G)$. This problem was introduced by Grünbaum [13], who proved that a graph with maximum degree 3 has an acyclic coloring with 4 colors. Similarly, an acyclic coloring of a graph $G$ was introduced in [13], as the proper coloring of $G$ not having any bicolored cycle. The minimum number of colors needed in an acyclic coloring of a graph $G$ is called the acyclic chromatic number of $G$, denoted by $a(G)$, where $a(G)\leq\chi_{s}(G).$ Both the star coloring and acyclic coloring problems were shown to be NP-complete by Albertson et al. [1] and Kostochka [19], respectively.
The star and acyclic coloring problems are also related to the proper coloring $G^{2}$ (square) of a graph $G$, where $V(G^{2})=V(G)$ and any two vertices are adjacent in $G^{2}$ if and only if they are at most distance two apart in $G$. In that sense, every proper coloring of $G^{2}$ is a star and acyclic coloring of $G.$ We study a variation of these problems avoiding bicolored paths. We call a proper vertex coloring of a graph $G$ without a bicolored copy of $P_{k}$ a $P_{k}$-coloring of $G$ for $k\geq 3$. The minimum number of colors needed for a $P_{k}$-coloring of $G$ is called $P_{k}$-chromatic number of $G,$ denoted by $s_{k}(G).$ We note that, for $k\geq 4$,
$\chi(G)\leq s_{k}(G)\leq s_{4}(G)=\chi_{s}(G)\leq\chi(G^{2})=s_{3}(G).$
Another relevant problem to $P_{k}$-coloring is nonrepetitive graph coloring. For a path $P=(v_{1},\dots,v_{k},v_{k+1},\dots,v_{2k}),$ we say $P$ a repetitively colored if the color of $v_{i}$ and $v_{k+i}$ are the same for each $1\leq i\leq k.$ A coloring of a graph $G$, where no path with even vertices is repetitively colored, is called a nonrepetitive coloring of $G.$ The (path-)nonrepetitive chromatic number, $\pi(G)$, is the minimum integer $k$ such that $G$ admits a nonrepetitive $k$-coloring. Since the seminal result by Alon et al. [2], nonrepetitive colorings of graphs have been widely studied. Wood [23] provides an extensive survey on this coloring. The $P_{k}$ coloring problem, $k\geq 4$, is closely related to this problem as
$s_{k}(G)\leq\chi_{s}(G)\leq\pi(G).$
The $P_{k}$-coloring problem was posed by Alon, McDiarmid, and Reed [3] who showed that for any graph $G$ with maximum degree $d,$ $a(G)=O(d^{\frac{4}{3}})$, and claimed that a similar upper bound can be shown for $P_{k}$-chromatic number. This coloring parameter is studied for paths of even order in [11] and later for all paths in [14, 18].
In this paper, we consider the $P_{k}$-coloring of 2-dimensional grid graphs, $P_{m}\square P_{n}$. The graph products have been studied widely on the problems listed above. In [12], Fertin, Godard and Raspaud provide both lower and upper bound for the acyclic chromatic number of $d$-dimensional grid, that is product of $d$ paths. In [15], Jamison and Matthews find the acyclic chromatic number of $P_{m}\square C_{n}$ (cylinder) and $C_{m}\square C_{n}$ (tori) as at most 4 and 5, respectively. Borodin [7] showed that every planar graph has an acyclic 5-coloring. Alon, Mohar, and Sanders [4] showed that each projective plane has an acyclic 7-coloring. Acyclic coloring of products of graphs are extensively studied also in [17, 16]. In all these results, we observe that the difference between the chromatic number and acyclic chromatic number is a constant value, independent from the length of the cycles that are avoided to be bicolored. Similarly, the most recent result on nonrepetitive coloring of grids is studied in [21], showing that $5\leq\pi(P_{n}\square P_{n})\leq 12$, for sufficiently large $n$. There are studies showing constant (upper and lower, resp.) bounds on the nonrepetitive chromatic number for graph families, such as paths (3,3) [22], cycles (3,4) [8], outerplanar graphs (7, 12) [5, 6, 20] and planar graphs (11,768) [10, 9]. As in these examples, our result in Theorem 1 shows that the $P_{k}$-chromatic number of the grid increases by a constant with respect to its chromatic number, independent of the value of $k$.
Theorem 1.
For any $k\geq 5$ and $m,n\geq k-2,$ $s_{k}(P_{m}\square P_{n})=4$.
In [18], Kırtışoğlu and the second author showed that $s_{k}(P_{k-3}\square P_{n})=3$ for all $k\geq 5$ and $n\geq 1,$ by providing the following colorings for the case $k=5,6,$ which can be generalized to all $k\geq 5.$
$\begin{matrix}1&2&3&1&2&3&...\\ 2&3&1&2&3&1&...\\ \end{matrix}\qquad\qquad\qquad\begin{matrix}1&2&3&1&2&3&\dots\\ 2&3&1&2&3&1&\dots\\ 1&2&3&1&2&3$$\dots\\ \end{matrix}$
This coloring pattern having columns with alternating colors from $(1,2),$ $(2,3),$ $(3,1),$ respectively, yields a valid 3-coloring for any $k\geq 6$ and $n\geq 1$, showing $s_{k}(P_{k-3}\square P_{n})=3$. Note that in such colorings, a bicolored $P_{k}$ has to have at least k2k-2 vertices in the same column, thus cannot be found in $P_{k-3}\square P_{n}$ colored according to the pattern above. In [18], it is also observed that for $k=5,6$, $s_{k}(P_{k-2}\square P_{n})=4$ for all $n\geq k-2$ and this is conjectured to hold for all $k.$ By the monotonicity of $s_{k}(G)$ with respect to $k$, this result implies, for any $k\geq 5$ and $m,n\geq k-2,$ $s_{k}(P_{m}\square P_{n})\leq 4$. With Theorem 1, we confirm this conjecture showing that there is no proper 3-coloring of $P_{m}\square P_{n}$ avoiding a bicolored $P_{k}$, for $m,n\geq k-2$. In the following section, we provide the proof of Theorem 1.
2. Main Result
We call a maximal connected subgraph induced by vertices having only two colors a bicolored component. To prove Theorem 1, we analyze bicolored components containing vertices from one of the sides of the grid. These components belong to one of the groups below:
  • peripheral bicolored component: a component that has vertices in two opposite sides of the grid, i.e., top and bottom sides, or left and right sides.
  • partial bicolored component: a component that is not peripheral, but has vertices on at least one of the sides of the grid.
In the rest, we assume that the sides of the grid that a partial component may intersect are top and left sides, since remaining cases are symmetric. We categorize each partial bicolored component $C$ as:
  • Type 1: if the vertices of $C$ are only on the top side,
  • Type 2: if the vertices of $C$ are on the top and left sides.
We see examples of Type 1 and Type 2 partial bicolored (as red-blue colored) components in Figure 1 (a), and in Figure 1 (b),(c), respectively. The following definitions are associated with a (partial or peripheral) bicolored component $C$.
Definition 1.
When considering a (bicolored) component $C$ as a planar subgraph in the grid drawing, the boundary of C is a walk traversing the outer face of $C$, starting at any vertex, in clockwise direction.
For example, in Figure 1 (c), the boundary of $C$ is the walk $(s,a,b,a,d,e,t,e,d,a,s).$ In Figure 1 (b), the boundary of $C$ is the walk $(s,a,b,c,d,c,t,c,e,a,s).$ In the rest, we make use of the term outside of the grid in the following sense.
Definition 2.
When walking clockwise on the boundary of $C$, if one’s left hand stays outside the grid along some segment $S$ of that walk, we say $S$ is outside the grid.
For example, in Figure 1 (c), the segment $(t,e,d,a,s)$ of the boundary $(s,a,b,a,d,e,t,e,d,a,s)$ is outside the grid.
Definition 3.
A partial walk $B_{C}$ is a maximal subwalk on the boundary of a (bicolored) component $C$ without a segment outside of the grid. In addition, if $C$ is a partial bicolored component, we let the starting vertex of $B_{C}$ be the rightmost vertex of $C$ on the top side of the grid.
By this definition, it is possible to have more than one partial walk on the boundary of $C$ only if $C$ is a peripheral bicolored component. Some examples for the partial walks are shown in Figure 1 (a), (b) and (c), where $B_{C}$ is described by the vertex sequence $(s,a,b,a,c,a,d,a,s)$, $(s,a,b,c,d,c,t)$, $(s,a,b,a,d,e,t)$, respectively. In Figure 1 (a), the boundary of $C$ happens to be the same as $B_{C}.$ In Figure 1 (d), for example, there are two red-blue colored peripheral components, one containing $z_{1}$, call it $C^{\prime}$, and another containing $w_{1}$, call it $C^{\prime\prime}.$ In $C^{\prime}$, there is only one partial walk $(y_{5},z_{5},z_{4},z_{3},y_{3},z_{3},z_{2},z_{1},y_{1}).$ In $C^{\prime\prime}$, there are several partial walks, such as $(w_{1},w_{2},x_{2},w_{2},w_{3},w_{4},x_{4},w_{4},w_{5})$, $(w_{5},w_{4},u_{4})$, $(u_{4},w_{4},w_{3},w_{2},u_{2})$ and $(u_{2},w_{2},w_{1}).$
In the rest, for simplicity, we write 3-coloring to refer to a proper 3-coloring of $V(P_{k-2}\square P_{k-2})$, $k\geq 5$, using colors red, blue and green, unless indicated otherwise. With Lemma 2, we make a generalization about the boundary structure of bicolored components.
Lemma 2.
For any $k\geq 5$, let $C$ be a (partial or peripheral) bicolored component in a 3-coloring $c:V(P_{k-2}\square P_{k-2})\to\{c_{1},c_{2},c_{3}\},$ and label the vertices along a partial walk $B_{C}$ as $v_{1},v_{2},\dots,v_{r}$. Then, the following hold:
  1. (1)
    $r\geq 3$.
  2. (2)
    The angles between the edges $v_{1}v_{2}$ and $v_{2}v_{3}$ and between $v_{r}v_{r-1}$ and $v_{r-1}v_{r-2}$ are $90^{\circ}$.
  3. (3)
    The angle between the edges $v_{i}v_{i+1}$ and $v_{i+1}v_{i+2}$, $i\leq r-2$, along $B_{C}$ is $90^{\circ}$ if and only if $i$ is odd.
  4. (4)
    $r$ is an odd integer.
  5. (5)
    Let $c(v_{1})=c_{1}$ and $C$ have colors $c_{1}$ and $c_{2}.$ Then, there is a $c_{1}$-$c_{3}$ colored connected subgraph containing the vertices $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}$ is the $c_{3}$-colored vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$
[…]
[scale=0.3]bndry_combined_four.jpg
Figure 1. In (a)-(c), examples of bicolored components show the edges of $B_{C}$ (bold edges) and the neighboring edges of another bicolored component (dashed edges). In (d), a coloring of the grid shows examples of peripheral bicolored components.
Proof.
For simplicity, let us consider the 3-coloring described above, using colors red, blue and green and let $C$ be colored with red and blue. Throughout the proof, we often make use of the fact that whenever a vertex on the partial walk $B_{C}$ has a neighbor outside $C$, this neighbor has color green.
(1) Since $C$ is bicolored, it has at least one edge. $C$ cannot be only a single edge, otherwise there are green colored adjacent vertices. For example, in Figure 1 (d), if $y_{1}z_{1}$ is the only edge in its bicolored component, then $y_{2}$ and $z_{2}$ would be green. Thus, $C$ has at least two edges, the smallest case for $C$ having only the two edges incident to the top left corner vertex. Thus, $r\geq 3.$
(2) Let the angle between $v_{1}v_{2}$ and $v_{2}v_{3}$ be $\theta$ and assume that $v_{1}v_{2}$ is a vertical edge, as in Figure 1 (a).
If $C$ is a partial bicolored component, recall that, $v_{1}$ is not part of the right side of the grid. Thus, in case $\theta$ has one of the values $180^{\circ}$, $270^{\circ}$ or $360^{\circ}$, by definition of $B_{C}$, the neighbors of $v_{1}$ and $v_{2}$ to their right are colored green and are adjacent, a contradiction. For example, in Figure 1 (a), if $B_{C}$ were $(s,a,c,a,d,a,s)$, then $b$ had color green, a contradiction with proper coloring of the grid.
If $C$ is a peripheral bicolored component, assume that $v_{1}$ is on the top side of the grid and $\theta\neq 90^{\circ}$. If $v_{1}$ is not on the top right corner of the grid, then same arguments above hold. Otherwise, $v_{1}v_{2}$ is an edge both vertical and an edge on the right side of the grid. However, this is not possible, as by our assumption, $B_{C}$ is a clockwise walk and does not contain a segment outside the grid. For example, in Figure 1 (d), the bicolored component containing $z_{1}$ and $z_{2}$, call it $Q$, has a partial walk as $(y_{5},z_{5},\dots,z_{2},z_{1},y_{1}).$ In this example, the segment $z_{1}z_{2}$ is a segment of the boundary of $Q$ outside of the grid. Hence, $\theta=90^{\circ}.$
If $v_{1}v_{2}$ is a horizontal edge as in Figure 1 (b) and (c), similar arguments as above show that $\theta=90^{\circ}.$ By symmetry, the angle between the edges $v_{r-2}v_{r-1}$ and $v_{r-1}v_{r}$ cannot be other than $90^{\circ}$.
(3) To clarify the definition of the angles along $B_{C}$, we first present some examples. In Figure 1 (a), the angles between the consecutive edges along $B_{C}$ are alternating between $90^{\circ}$ and $360^{\circ},$ whereas in Figure 1 (c), these angles from $s$ to $t$ are, respectively, $90^{\circ},$ $360^{\circ}$, $90^{\circ}$, $180^{\circ}$ and $90^{\circ}$. In Figure 1 (d), for the red-blue colored component with $z_{1}$ in it, these angles along the only partial walk (starting at $y_{5}$ ending at $y_{1}$) follow as $90^{\circ},$ $180^{\circ}$, $90^{\circ}$, $360^{\circ}$, $90^{\circ}$, $180^{\circ}$, $90^{\circ}$.
Let $w,x,y,z$ represent the vertices $v_{i},v_{i+1},v_{i+2},v_{i+3}$, respectively, for some $i.$ Call the angle between the edge pairs $(wx,xy)$ and $(xy,yz)$, $\beta$ and $\gamma,$ respectively. In the following, we discuss invalid values for $(\beta,\gamma)$, omitting symmetric cases and letting $w$ have color red. All cases are illustrated in Figure 2.
[…]
[width=0.8]Case2_invalid_combin.jpg
Figure 2. Invalid cases listing values of $(\beta,\gamma)$.
  • $(\beta,\gamma)=(180^{\circ},180^{\circ}),(270^{\circ},180^{\circ}),(270^{\circ},270^{\circ})$: As $B_{C}$ has no segment outside the grid, $x$ and $y$ have neighbors outside $C$ that are adjacent and green.
  • $(\beta,\gamma)=(360^{\circ},180^{\circ}),(360^{\circ},270^{\circ})$: In each of these cases, we use the fact that $x$ is not on any side of the grid, as in this ordering, $x$ cannot be $v_{1}$ or $v_{r}$ (endvertices of $B_{C}$). Thus, one of the $C_{4}$’s containing the edge $xy$, shown in Figure 2, has two adjacent vertices that are not in $C$, hence green, a contradiction.
  • $(\beta,\gamma)=(360^{\circ},360^{\circ})$: This case is not possible as it implies that $C$ consists of a single edge contradicting (1).
Thus, we observe that either $\beta$ or $\gamma$ is $90^{\circ}$ as listed in Figure 3, where the vertices marked with * are implied to be in the bicolored component $C$ to have a proper coloring. In Figure 3 the case $(\beta,\gamma)=(90^{\circ},90^{\circ})$ gives a contradiction since the segment of $B_{C}$ should not be $(w,x,y,z)$ but $(w,z)$. Thus, (3) holds.
[…]
[scale=0.3]Case2_valid.jpg
Figure 3. The only valid configurations for $(\beta,\gamma)$ (omitting symmetric cases), except the leftmost case. The vertices marked with * are part of $C.$
(4) This is implied by (2) and (3).
(5) Let $u_{(i+1)/2}$ be the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\},$ for each odd $1\leq i\leq r-2$. Note that $u_{(i+1)/2}$ is not on $B_{C}$, since two consecutive degrees along $B_{C}$ cannot be both $90^{\circ}$. Also, vertices $u_{(i+1)/2}$ are not necessarily distinct, for example, in Figure 1 (c), $u_{2}=u_{3}=c$. As given in the statement, let $c_{1}$ be the color of $v_{1}$ and let $C$ be $c_{1}$-$c_{2}$ colored. The edges $v_{i}u_{(i+1)/2}$ and $v_{i+2}u_{(i+1)/2}$ have colors $c_{3}$ and $c_{1}$, for each odd $i$, $1\leq i\leq r-2.$ So, we obtain a $c_{1}$-$c_{3}$ colored connected subgraph induced by the vertices $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2.$ Some examples of this bicolored subgraph neighboring $C$ are shown in Figure 1 with dashed edges. ∎
Lemma 3 below is used to show that Theorem 1 holds in case there is a peripheral bicolored component in a 3-coloring of $P_{k-2}\square P_{k-2}$, $k\geq 5$.
Lemma 3.
For any $k\geq 5$, if a 3-coloring of $P_{k-2}\square P_{k-2}$ has a peripheral bicolored component, then there is a bicolored $P_{k}$.
Proof.
If the 3-coloring of $P_{k-2}\square P_{k-2}$ has a peripheral bicolored component, then let $P$ be a maximal bicolored path connecting two opposite sides, w.l.o.g. top and bottom sides, of the grid. We label the columns of the grid as $C_{1},\dots,$ $C_{k-2}$ from left to right. Let $C_{j}$ be the leftmost column containing vertices from $P$. If they exist, we label the vertex set in $C_{j-1},C_{j},C_{j+1},C_{j+2}$ as $v_{s},w_{s},x_{s},y_{s},$ $1\leq s\leq k-2$, respectively, $s$ indicating the row index with $s=1$ being the top row. Note that, it is possible that some of these columns do not exist depending on whether $P$ has edges on the right or left side of the grid. We present below a case analysis considering these possibilities as well and find a bicolored $P_{k}$ in each case. By the maximality of $P$, there can be at most one adjacent pair of vertices from different columns (jump across columns) in P, otherwise we have a bicolored $P_{k}.$ In that sense, $P$ has vertices in at most two columns. Assume, w.l.o.g., that $P$ is a red-blue colored path and $w_{1}$ (the leftmost top vertex of $P$) has color red in the rest of the proof.
Case 1. |V(P)|=k2.|V(P)|=k-2.
The only possibility of this case is that $P$ contains only vertices from a single column (only $C_{j}$) having exactly k2k-2 vertices, i.e. $V(P)=\{w_{1},w_{2},\dots,w_{k-2}\}.$ Assuming that $C_{j+1}$ exists, $x_{1}$ has color green by the maximality of $P$. Since $w_{2}$ has color blue, $x_{2}$ has color red for a proper coloring. Also, the remaining vertices on $C_{j+1}$ have only red and green colors by maximality of $P$. Hence, the vertices on $C_{j+1}$ together with $w_{1}$ induce a red-green colored $P_{k-1}$. If $C_{j+1}$ does not exist and $P$ is the right side of the grid, then the same arguments yield a red-green colored $P_{k-1}$ similarly using the vertices on $C_{j-1}$ and $w_{1}.$ So, we consider this case in the following.
Case 2. |V(P)|=k1.|V(P)|=k-1.
Let $P$ have a jump from $C_{j}$ to $C_{j+1}$ at row $i,$ 1<i<k2.1<i<k-2. We consider the case i=1,k2i=1,k-2 at the end of this case. We only consider the jump of $P$ from left ($C_{j}$) to right ($C_{j+1}$), since considering a jump from right to left is covered by symmetry. Thus, we have $V(P)=\{w_{1},\dots,w_{i},x_{i},\dots,x_{k-2}\}.$ In the following, we discuss possible colorings depending on the parity of $i$ and $k$, listed in Figure 4, respectively. To find a bicolored $P_{k}$, we repetitively use the following remark.
Remark 4.
Given the description of $P$ above and assuming that $w_{1}$ is colored red, we have that
  • the paths $v_{1},\dots,v_{i}$ (if they exist) and $x_{1},\dots,x_{i-1}$ are red-green colored paths,
  • the paths $w_{i+1},\dots,w_{k-2}$ and $y_{i},\dots,y_{k-2}$ (if exists) are $\alpha$-green colored paths, where $\alpha$ is the color of $x_{k-2}$.
Moreover, the endvertices of these paths $v_{1}$, $x_{1}$, $w_{k-2}$, $y_{k-2}$ have color green whereas the color of $v_{i},x_{i-1},w_{i+1},y_{i}$ depend on the parity of $i$ and $k.$
Proof.
The color of $x_{1}$ is green by maximality of $P$ and $x_{2}$ is red in a proper coloring, because $w_{2}$ is blue by our assumption. For each $3\leq i^{\prime}\leq i-1$, one can find the color of $x_{i^{\prime}}$ inductively: The color of $x_{i^{\prime}}$ is green if $i^{\prime}$ is odd, because $x_{i^{\prime}-1}$ has color red by the inductive hypothesis and $w_{i^{\prime}}$ has color red because of our assumption that $w_{1}$ is red. Note that if $x_{i^{\prime}}$ has color blue, we obtain a longer bicolored path by the union of $P$, $x_{i^{\prime}-1}$ and $x_{i^{\prime}}$, a contradiction. The color of $x_{i^{\prime}}$ is red if $i^{\prime}$ is even, because $x_{i^{\prime}-1}$ has color green by the inductive hypothesis and $w_{i^{\prime}}$ has color blue because of our assumption that $w_{1}$ is red. Also, if $C_{j-1}$ exists, that is $j\not=1$, we observe that $v_{1}$ is green by maximality of $P$ and $v_{1},v_{2},\dots,v_{i}$ is a red-green path by similar reasoning as above.
Let $\alpha$ be the color of $x_{k-2},$ where $\alpha$ is either red or blue, depending on the modularity of $k.$ Let $\bar{\alpha}$ be the color on $P$ that is different than $\alpha,$ that is $\bar{\alpha}$ is blue if $\alpha$ is red and vice versa. The color of $w_{k-2}$ is green by maximality of $P$ and $w_{k-3}$ has color $\alpha$ in a proper coloring, since $x_{k-3}$ has color $\bar{\alpha}$. For each $4\leq i^{\prime}\leq k-i-1$, one can find the color of $w_{k-i^{\prime}}$ inductively: The color of $w_{k-i^{\prime}}$ is green if $i^{\prime}$ is even, since $w_{k-(i^{\prime}-1)}$ has color $\alpha$ by the inductive hypothesis and $x_{k-i^{\prime}}$ has color $\alpha$ as well. Note that if $w_{k-i^{\prime}}$ has color $\bar{\alpha}$, we obtain a longer bicolored path by the union of $P$, $w_{k-i^{\prime}}$ and $w_{k-(i^{\prime}-1)}$, a contradiction. Similarly, the color of $w_{k-i^{\prime}}$ is $\alpha$ if $i^{\prime}$ is odd, because $w_{k-(i^{\prime}-1)}$ is green by the inductive hypothesis and $x_{k-i^{\prime}}$ has color $\bar{\alpha}$. If $C_{j+2}$ exists, that is $j\leq k-4$, $y_{k-2}$ is green by maximality of $P$ and by similar reasoning we obtain that $y_{i+1},\dots,y_{k-2}$ is a $\alpha$-green path.
As the color of $v_{i},x_{i-1},w_{i+1},y_{i}$ depend on the parity of $i$ and $k,$ the discussion on determining the color of these vertices is continued in the case analysis below.
[…]
[scale=0.4]Case1_maximality.jpg
Figure 4. The neighboring columns of maximal red-blue colored $P$ (bold edges) and possible colorings: (a) $i\not\equiv k$ (mod 2) and $i$, even, (b) $i\equiv k$ (mod 2) and $i$, even, (c) $i\not\equiv k$ (mod 2) and i, odd, (d) $i\equiv k$ (mod 2) and $i$, odd.
Case 2.1. $i\not\equiv k$ (mod 2) and $i$, even.
If $j\leq k-4$, then $C_{j+2}$ exists and by Remark 4, there is a red-green $P_{2i+1}$ induced by $\{x_{1},\dots,x_{i},w_{1}$, $v_{1},\dots,v_{i}\}$ and a blue-green $P_{2(k-i-1)+1}$ induced by $\{w_{i},\dots,w_{k-2},x_{k-2},y_{i},\dots,y_{k-2}\}$, as shown in Figure 4 (a). One of these bicolored paths has at least $k$ vertices, depending on the value of $i$.
If j=k3j=k-3, then $C_{j+2}$ does not exist. In this case, it is not guaranteed to have a bicolored $P_{k}$ by the neighbors of $V(P)$ and itself. So, we use Lemma 2 to show the existence of a bicolored $P_{k}$. Let $C$ be the bicolored red-blue component containing $P.$ If $C$ contains any vertex from the top or bottom sides outside $P,$ then $C$ contains a red-blue $P_{k}.$ If $C$ contains any vertex from the left side outside $P,$ then the red-blue path in $C$ connecting $P$ to the left side together with $P$ yield a bicolored path with at least $k$ vertices. Thus, we assume that $C$ is not connected to any side of the grid (other than the right side) outside $P$. This implies that there exists a partial walk of $C$ from $x_{k-2}$ to $w_{1}$. However, Lemma 2 yields that $x_{k-2}$ and $w_{1}$ have the same color, a contradiction.
Case 2.2. $i\equiv k$ (mod 2) and $i$, even.
[…]
Case 2.3. $i\not\equiv k$ (mod 2) and i, odd.
If $j\not=1$, by Remark 4, the union of the two red-green paths $v_{1},v_{2},\dots,v_{i}$ and $w_{i},\dots,w_{k-2}$, together with $x_{k-2}$ induce a bicolored $P_{k}.$ If $j=1$, then $C_{j+2}$ exists. By Remark 4, $x_{1},\dots,x_{i-1}$ and $y_{i},\dots,y_{k-2}$ are red-green paths. If $y_{i-1}$ (black vertex in Figure 4 (c)) has color blue, then there is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},y_{i-1},y_{i},x_{i},$ $\dots,x_{k-2}\}.$ Otherwise, $y_{i-1}$ has color green and there is a red-green $P_{k}$ induced by $\{w_{1},x_{1},\dots x_{i-1},$ $y_{i-1},\dots,y_{k-2},x_{k-2}\}.$
Case 2.4. $i\equiv k$ (mod 2) and $i$, odd.
There is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},x_{i},w_{i},w_{i+1},x_{i+1},\dots,x_{k+2}\}.$
As a special case, if $i=1$, then $P$ crosses between columns at the top row having vertex set $\{w_{1},x_{1},x_{2},\dots,x_{k-2}\}$. By Remark 4, $w_{2},\dots,w_{k-2}$ induce a red-green path and $k$ must be even, since $w_{k-2}$ is green by maximality of $P$. If $C_{j-1}$ exists, $v_{1}$ can be either green or blue in a proper coloring, inducing either a red-green $P_{k}$ together with $w_{1},\dots,w_{k-2},x_{k-2}$ or a red-blue $P_{k}$ together with $P$, respectively. In case, $j=1$, then $C_{j+2}$ exists and the path $y_{2},\dots,y_{k-2}$ is a red-green colored path by Remark 4. The red-green colored paths on $C_{j}$ and $C_{j+2}$ together $x_{k-2}$ yield a red-green $P_{k}.$ The same proof holds also for i=k2i=k-2 by symmetry.
Proof of Theorem 1: We know that $s_{k}(P_{m}\square P_{n})\leq 4$ as $s_{5}(P_{m}\square P_{n})=4$, for any $m,n\geq 3$, $k\geq 5$ ([18]). To prove that $s_{k}(P_{m}\square P_{n})\geq 4$, it suffices to show that $s_{k}(P_{k-2}\square P_{k-2})\geq 4.$ We consider any 3-coloring of $P_{k-2}\square P_{k-2}$ for $k\geq 5$ and show in Claim 5 that it has a peripheral bicolored component. Hence, there is a bicolored $P_{k}$ by Lemma 3.
Let $P_{k-2}\square P_{k-2}$ be colored with colors $\{c_{1},c_{2},c_{3}\}$ and assume that such a 3-coloring does not have a peripheral bicolored component. Let $C$ be a partial, $c_{1}$-$c_{2}$ colored component containing vertices from the top side and possibly left side of the grid. Let $B_{C}=(v_{1},v_{2},\dots,v_{r})$ be a partial walk of $C$. We need the following definitions associated with $C$:
  • $D_{C}$: By Lemma 2, there is a $c_{1}$-$c_{3}$ colored connected subgraph, $c_{1}$ being the color of $v_{1}$, induced by $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}\not\in C$ is the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$ Let $D_{C}$ be the bicolored component containing this $c_{1}$-$c_{3}$ colored connected subgraph.
  • $\mathcal{B}_{C}$: Let this be the collection of $C_{4}$’s induced by $\{u_{(i+1)/2},v_{i},v_{i+1},v_{i+2}\}$, for odd $i$, $1\leq i\leq r-2$. As $r\geq 3$, $\mathcal{B}_{C}\neq\emptyset$.
  • $R_{C}$: This is the region of the grid enclosed by $B_{C}$ and the walk along the four sides of the grid in clockwise direction starting at $v_{1}$ ending at $v_{r}$. In other words, it is the region to the left when walking on $B_{C}$ from $v_{1}$ to $v_{r}$ that is enclosed by the four sides of the grid.
Claim 5.
When $C$ is replaced with $D_{C}$ iteratively, $D_{C}$ will be a peripheral bicolored component after a finite number of iterations.
Proof.
By definition and by Lemma 2, at each iteration, $D_{C}$ contains the vertices $v_{1}$ and $v_{r}$ of $B_{C}=(v_{1},\dots,v_{r})$, where $v_{1}$ is on the top side and $v_{r}$ is either on the top side or left side of the grid. Let $B_{C}^{\prime}=(v_{1}^{\prime},\dots,v_{r}^{\prime})$ be the partial walk of $D_{C}$. Assuming $D_{C}$ is not peripheral, possible scenarios at each iteration are:
  1. (1)
    $C$ and $D_{C}$ have some vertices only from the top side and no other side of the grid. Thus, $v_{1}$, $v_{1}^{\prime}$, $v_{r}$, $v_{r}^{\prime}$ are on the top side of the grid.
  2. (2)
    $C$ has some vertices only from the top side and no other side of the grid, whereas $D_{C}$ has vertices from the top and one of the right or left sides of the grid. In case, $D_{C}$ has some vertex on the right side, we flip the grid to exchange the right and left sides of the grid so that $D_{C}$ satisfies our assumptions. Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side, $v_{r}$ is on the top side, $v_{r}^{\prime}$ is on the left side.
  3. (3)
    Both $C$ and $D_{C}$ have some vertices from the top and left side of the grid. Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side of the grid, $v_{r}$, $v_{r}^{\prime}$ are on the left side of the grid.
In these cases, the grid is not replaced with any of its automorphisms and the position of the vertices on the grid do not change at any iteration, except in (2). Note that, an iteration in (2) may occur at most once. In all iterations, except possibly one in case (2), we observe that $v_{1}^{\prime}$ is to the right of $v_{1}$ on the top side of the grid or $v_{1}=v_{1}^{\prime}$. Similarly, in all iterations in (1), $v_{r}^{\prime}$ is to the left of $v_{r}$ on the top side of the grid or $v_{r}=v_{r}^{\prime}$, and in (3), $v_{r}^{\prime}$ is below $v_{r}$ on the left side of the grid or $v_{r}=v_{r}^{\prime}$.
Note that, after an iteration, the squares enclosed by the $C_{4}$’s in $\mathcal{B}_{C}$ are missing in $R_{C}$. As $\mathcal{B}_{C}\neq\emptyset$, $R_{C}$ becomes smaller after each iteration in cases (1)-(3). Thus, as long as $C$ is not peripheral, $v_{1}=v_{1}^{\prime}$ or $v_{r}=v_{r}^{\prime}$ can happen at most once for each vertex, as there are exactly two $C_{4}$’s that contain $v_{1}$ (same for $v_{r}$) and no $C_{4}$ belongs to $R_{C}$ in more than one iteration. In that sense, each vertex plays the role of $v_{1}$ or $v_{r}$ with multiplicity at most two throughout all iterations. Consequently, at each consecutive pair of iterations, the position of $v_{1}$ approaches the top right corner by at least one edge and the position of $v_{r}$, approaches the top left or the bottom left corner of the grid by at least one edge, depending on the case. These observations show that after some finite number of iterations, none of the cases (1)-(3) will hold and $D_{C}$ becomes a peripheral bicolored component. ∎
Acknowledgements
The authors would like to thank the referees for their valuable comments that improved the quality of this manuscript.

### v5

@@
 Abstract.
-The star chromatic number on a graph is the minimum number of colors in a proper vertex coloring forbidding any $P_{4}$ with two colors (bicolored).
+The star chromatic number of a graph is the minimum number of colors in a proper vertex coloring forbidding any $P_{4}$ with two colors (bicolored).
 This problem was introduced by Grünbaum (1973) together with the acyclic coloring of graphs, where bicolored cycles are avoided.
@@
 1. Introduction
-For a graph $G$, $\chi(G)$ denotes the chromatic number of $G$, that is the minimum number of colors needed to color $V(G)$ without a monochromatic edge.
+For a graph $G$, $\chi(G)$ denotes the chromatic number of $G$, that is, the minimum number of colors needed to color $V(G)$ without a monochromatic edge.
 The star coloring problem on a graph $G$ asks to find the minimum number of colors in a proper coloring forbidding a bicolored (2-colored) $P_{4},$ called the star-chromatic number $\chi_{s}(G)$.
@@
 [1] and Kostochka [19], respectively.
-The star and acyclic coloring problems are also related to the proper coloring $G^{2}$ (square) of a graph $G$, where $V(G^{2})=V(G)$ and any two vertices are adjacent in $G^{2}$ if and only if they are at most distance two apart in $G$.
+The star and acyclic coloring problems are also related to the proper coloring of $G^{2}$, the square of a graph $G$, where $V(G^{2})=V(G)$ and any two vertices are adjacent in $G^{2}$ if and only if they are at distance at most two in G.
 In that sense, every proper coloring of $G^{2}$ is a star and acyclic coloring of $G.$ We study a variation of these problems avoiding bicolored paths.
@@
 Another relevant problem to $P_{k}$-coloring is nonrepetitive graph coloring.
-For a path $P=(v_{1},\dots,v_{k},v_{k+1},\dots,v_{2k}),$ we say $P$ a repetitively colored if the color of $v_{i}$ and $v_{k+i}$ are the same for each $1\leq i\leq k.$ A coloring of a graph $G$, where no path with even vertices is repetitively colored, is called a nonrepetitive coloring of $G.$ The (path-)nonrepetitive chromatic number, $\pi(G)$, is the minimum integer $k$ such that $G$ admits a nonrepetitive $k$-coloring.
+For a path $P=(v_{1},\dots,v_{k},v_{k+1},\dots,v_{2k}),$ we say $P$ a repetitively colored if the color of $v_{i}$ and $v_{k+i}$ are the same for each $i$, $1\leq i\leq k.$ A coloring of a graph $G$, where no path with even vertices is repetitively colored, is called a nonrepetitive coloring of $G.$ The (path-)nonrepetitive chromatic number, $\pi(G)$, is the minimum integer $k$ such that $G$ admits a nonrepetitive $k$-coloring.
 Since the seminal result by Alon et al.
@@
 In this paper, we consider the $P_{k}$-coloring of 2-dimensional grid graphs, $P_{m}\square P_{n}$.
-The graph products have been studied widely on the problems listed above.
+Graph products have been studied widely for the problems listed above.
 In [12], Fertin, Godard and Raspaud provide both lower and upper bound for the acyclic chromatic number of $d$-dimensional grid, that is product of $d$ paths.
-In [15], Jamison and Matthews find the acyclic chromatic number of $P_{m}\square C_{n}$ (cylinder) and $C_{m}\square C_{n}$ (tori) as at most 4 and 5, respectively.
+In [15], Jamison and Matthews find the acyclic chromatic number of $P_{m}\square C_{n}$ (cylinder) and $C_{m}\square C_{n}$ (torus) as at most 4 and 5, respectively.
 Borodin [7] showed that every planar graph has an acyclic 5-coloring.
@@
 We call a maximal connected subgraph induced by vertices having only two colors a bicolored component.
+In particular, a bicolored component that has colors, say $c_{1}$ and $c_{2}$, is called a $c_{1}$-$c_{2}$ colored component.
 To prove Theorem 1, we analyze bicolored components containing vertices from one of the sides of the grid.
-These components belong to one of the groups below:
-  • peripheral bicolored component: a component that has vertices in two opposite sides of the grid, i.e., top and bottom sides, or left and right sides.
-  • partial bicolored component: a component that is not peripheral, but has vertices on at least one of the sides of the grid.
-In the rest, we assume that the sides of the grid that a partial component may intersect are top and left sides, since remaining cases are symmetric.
+These bicolored components belong to one of the groups below:
+  • peripheral bicolored component: a bicolored component that has vertices in two opposite sides of the grid, i.e., top and bottom sides, or left and right sides.
+  • partial bicolored component: a bicolored component that is not peripheral, but has vertices on at least one of the sides of the grid.
+In the rest, we assume that the sides of the grid that a partial bicolored component may intersect are top and left sides, since remaining cases are symmetric.
 We categorize each partial bicolored component $C$ as:
@@
   • Type 2: if the vertices of $C$ are on the top and left sides.
-We see examples of Type 1 and Type 2 partial bicolored (as red-blue colored) components in Figure 1 (a), and in Figure 1 (b),(c), respectively.
+We see examples of Type 1 and Type 2 partial bicolored components (as red-blue colored components) in Figure 1 (a), and in Figure 1 (b),(c), respectively.
 The following definitions are associated with a (partial or peripheral) bicolored component $C$.
 Definition 1.
-When considering a (bicolored) component $C$ as a planar subgraph in the grid drawing, the boundary of C is a walk traversing the outer face of $C$, starting at any vertex, in clockwise direction.
+When considering a bicolored component $C$ as a planar subgraph in the grid drawing, the boundary of C is a walk traversing the outer face of $C$, starting at any vertex, in clockwise direction.
 For example, in Figure 1 (c), the boundary of $C$ is the walk $(s,a,b,a,d,e,t,e,d,a,s).$ In Figure 1 (b), the boundary of $C$ is the walk $(s,a,b,c,d,c,t,c,e,a,s).$ In the rest, we make use of the term outside of the grid in the following sense.
@@
 When walking clockwise on the boundary of $C$, if one’s left hand stays outside the grid along some segment $S$ of that walk, we say $S$ is outside the grid.
+It is possible that a segment outside the grid is a single vertex.
 For example, in Figure 1 (c), the segment $(t,e,d,a,s)$ of the boundary $(s,a,b,a,d,e,t,e,d,a,s)$ is outside the grid.
 Definition 3.
-A partial walk $B_{C}$ is a maximal subwalk on the boundary of a (bicolored) component $C$ without a segment outside of the grid.
-In addition, if $C$ is a partial bicolored component, we let the starting vertex of $B_{C}$ be the rightmost vertex of $C$ on the top side of the grid.
-By this definition, it is possible to have more than one partial walk on the boundary of $C$ only if $C$ is a peripheral bicolored component.
+Given a fixed configuration of the grid, a partial walk $B_{C}$ is a maximal subwalk of the boundary of a bicolored component $C$ without any segment outside of the grid.
+In addition, if $C$ is a partial bicolored component of Type-1 or Type-2, we let the starting vertex of $B_{C}$ be the rightmost vertex of $C$ on the top side of the grid.
+To have a well-defined partial walk $B_{C}$, we assume that if $C$ is a partial bicolored component of Type-1 or Type-2, the boundary of $C$ is a walk in clockwise direction and the starting vertex of $B_{C}$ be the rightmost vertex of $C$ on the top side of the grid.
+Thus, on a fixed configuration of the grid, it is possible to have more than one partial walk on the boundary of $C$ only if $C$ is a peripheral bicolored component.
 Some examples for the partial walks are shown in Figure 1 (a), (b) and (c), where $B_{C}$ is described by the vertex sequence $(s,a,b,a,c,a,d,a,s)$, $(s,a,b,c,d,c,t)$, $(s,a,b,a,d,e,t)$, respectively.
-In Figure 1 (a), the boundary of $C$ happens to be the same as $B_{C}.$ In Figure 1 (d), for example, there are two red-blue colored peripheral components, one containing $z_{1}$, call it $C^{\prime}$, and another containing $w_{1}$, call it $C^{\prime\prime}.$ In $C^{\prime}$, there is only one partial walk $(y_{5},z_{5},z_{4},z_{3},y_{3},z_{3},z_{2},z_{1},y_{1}).$ In $C^{\prime\prime}$, there are several partial walks, such as $(w_{1},w_{2},x_{2},w_{2},w_{3},w_{4},x_{4},w_{4},w_{5})$, $(w_{5},w_{4},u_{4})$, $(u_{4},w_{4},w_{3},w_{2},u_{2})$ and $(u_{2},w_{2},w_{1}).$
+In Figure 1 (a), the boundary of $C$ happens to be the same as $B_{C}.$ In Figure 1 (d), the red-blue colored peripheral component containing $z_{1}$ has only one partial walk $(y_{6},y_{5},x_{5},y_{5},y_{4},y_{3},x_{3},y_{3},y_{2},y_{1},x_{1}).$ The green-blue colored peripheral component containing $w_{1}$, has several partial walks, for example $(v_{6},v_{5},u_{5})$ and $(w_{1},w_{2},x_{2},w_{2},w_{3},w_{4},x_{4},w_{4},w_{5},v_{5},v_{6}).$ The partial red-blue colored component containing $v_{1}$ has only one partial walk $(v_{1},v_{2},w_{2},v_{2},v_{3},v_{4},w_{4},$ $v_{4},v_{5},v_{4},u_{4}).$
 In the rest, for simplicity, we write 3-coloring to refer to a proper 3-coloring of $V(P_{k-2}\square P_{k-2})$, $k\geq 5$, using colors red, blue and green, unless indicated otherwise.
 With Lemma 2, we make a generalization about the boundary structure of bicolored components.
+To clarify the definition of the angles along $B_{C}$, we first present some examples.
+In Figure 1 (a), the angles between the consecutive edges along $B_{C}$ are alternating between $90^{\circ}$ and $360^{\circ},$ whereas in Figure 1 (c), these angles from $s$ to $t$ are, respectively, $90^{\circ},$ $360^{\circ}$, $90^{\circ}$, $180^{\circ}$ and $90^{\circ}$.
+In Figure 1 (d), for the red-blue colored component with $z_{1}$ in it, these angles along the only partial walk (starting at $y_{6}$ ending at $x_{1}$) follow as $90^{\circ},$ $360^{\circ}$, $90^{\circ}$, $180^{\circ}$, $90^{\circ}$, $360^{\circ}$, $90^{\circ},180^{\circ},90^{\circ}.$
 Lemma 2.
@@
 […]
-[scale=0.3]bndry_combined_four.jpg
+[scale=0.3]bndry_combined_four_new.jpg
 Figure 1.
@@
 If $v_{1}$ is not on the top right corner of the grid, then same arguments above hold.
-Otherwise, $v_{1}v_{2}$ is an edge both vertical and an edge on the right side of the grid.
+Otherwise, $v_{1}v_{2}$ is a vertical edge on the right side of the grid.
 However, this is not possible, as by our assumption, $B_{C}$ is a clockwise walk and does not contain a segment outside the grid.
-For example, in Figure 1 (d), the bicolored component containing $z_{1}$ and $z_{2}$, call it $Q$, has a partial walk as $(y_{5},z_{5},\dots,z_{2},z_{1},y_{1}).$ In this example, the segment $z_{1}z_{2}$ is a segment of the boundary of $Q$ outside of the grid.
 Hence, $\theta=90^{\circ}.$
 If $v_{1}v_{2}$ is a horizontal edge as in Figure 1 (b) and (c), similar arguments as above show that $\theta=90^{\circ}.$ By symmetry, the angle between the edges $v_{r-2}v_{r-1}$ and $v_{r-1}v_{r}$ cannot be other than $90^{\circ}$.
-(3) To clarify the definition of the angles along $B_{C}$, we first present some examples.
-In Figure 1 (a), the angles between the consecutive edges along $B_{C}$ are alternating between $90^{\circ}$ and $360^{\circ},$ whereas in Figure 1 (c), these angles from $s$ to $t$ are, respectively, $90^{\circ},$ $360^{\circ}$, $90^{\circ}$, $180^{\circ}$ and $90^{\circ}$.
-In Figure 1 (d), for the red-blue colored component with $z_{1}$ in it, these angles along the only partial walk (starting at $y_{5}$ ending at $y_{1}$) follow as $90^{\circ},$ $180^{\circ}$, $90^{\circ}$, $360^{\circ}$, $90^{\circ}$, $180^{\circ}$, $90^{\circ}$.
-Let $w,x,y,z$ represent the vertices $v_{i},v_{i+1},v_{i+2},v_{i+3}$, respectively, for some $i.$ Call the angle between the edge pairs $(wx,xy)$ and $(xy,yz)$, $\beta$ and $\gamma,$ respectively.
-In the following, we discuss invalid values for $(\beta,\gamma)$, omitting symmetric cases and letting $w$ have color red.
+(3) Let $w,x,y,z$ represent the vertices $v_{i},v_{i+1},v_{i+2},v_{i+3}$, respectively, for some $i.$ Call the angle between the edge pairs $(wx,xy)$ and $(xy,yz)$, $\beta$ and $\gamma,$ respectively.
+In the following, we discuss invalid values for $(\beta,\gamma)$, letting $w$ have color red.
+We omit the symmetric cases, where the listed pairs of angles appear in reverse order.
 All cases are illustrated in Figure 2.
@@
 In Figure 3 the case $(\beta,\gamma)=(90^{\circ},90^{\circ})$ gives a contradiction since the segment of $B_{C}$ should not be $(w,x,y,z)$ but $(w,z)$.
-Thus, (3) holds.
+Thus, (3) follows from (2).
 […]
 [scale=0.3]Case2_valid.jpg
-Figure 3. The only valid configurations for $(\beta,\gamma)$ (omitting symmetric cases), except the leftmost case. The vertices marked with * are part of $C.$
+Figure 3.
+The only valid configurations for $(\beta,\gamma)$, except the leftmost case.
+The vertices marked with * are part of $C.$ (Symmetric cases are omitted, where the listed pairs of angles appear in reverse order.)
 (4) This is implied by (2) and (3).
-(5) Let $u_{(i+1)/2}$ be the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\},$ for each odd $1\leq i\leq r-2$.
+(5) Let $u_{(i+1)/2}$ be the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\},$ for each odd $i$, $1\leq i\leq r-2$.
 Note that $u_{(i+1)/2}$ is not on $B_{C}$, since two consecutive degrees along $B_{C}$ cannot be both $90^{\circ}$.
@@
 We present below a case analysis considering these possibilities as well and find a bicolored $P_{k}$ in each case.
-By the maximality of $P$, there can be at most one adjacent pair of vertices from different columns (jump across columns) in P, otherwise we have a bicolored $P_{k}.$ In that sense, $P$ has vertices in at most two columns.
-Assume, w.l.o.g., that $P$ is a red-blue colored path and $w_{1}$ (the leftmost top vertex of $P$) has color red in the rest of the proof.
+There can be at most one adjacent pair of vertices from different columns (jump across columns) in P, otherwise we have a bicolored $P_{k}.$ Thus, $P$ has vertices in at most two columns.
+Assume, w.l.o.g., that $P$ is a red-blue colored path and $w_{1}$ has color red in the rest of the proof.
 Case 1. |V(P)|=k2.|V(P)|=k-2.
-The only possibility of this case is that $P$ contains only vertices from a single column (only $C_{j}$) having exactly k2k-2 vertices, i.e.
+The only possibility of this case is that $P$ contains only vertices from a single column (only $C_{j}$), i.e.
 $V(P)=\{w_{1},w_{2},\dots,w_{k-2}\}.$ Assuming that $C_{j+1}$ exists, $x_{1}$ has color green by the maximality of $P$.
@@
 Hence, the vertices on $C_{j+1}$ together with $w_{1}$ induce a red-green colored $P_{k-1}$.
-If $C_{j+1}$ does not exist and $P$ is the right side of the grid, then the same arguments yield a red-green colored $P_{k-1}$ similarly using the vertices on $C_{j-1}$ and $w_{1}.$ So, we consider this case in the following.
+If $C_{j+1}$ does not exist and $P$ is the right side of the grid, then the same arguments yield a red-green colored $P_{k-1}$ similarly using the vertices on $C_{j-1}$ and $w_{1}.$ This contradicts the maximality of $P.$
 Case 2. |V(P)|=k1.|V(P)|=k-1.
-Let $P$ have a jump from $C_{j}$ to $C_{j+1}$ at row $i,$ 1<i<k2.1<i<k-2.
-We consider the case i=1,k2i=1,k-2 at the end of this case.
-We only consider the jump of $P$ from left ($C_{j}$) to right ($C_{j+1}$), since considering a jump from right to left is covered by symmetry.
-Thus, we have $V(P)=\{w_{1},\dots,w_{i},x_{i},\dots,x_{k-2}\}.$ In the following, we discuss possible colorings depending on the parity of $i$ and $k$, listed in Figure 4, respectively.
-To find a bicolored $P_{k}$, we repetitively use the following remark.
-Remark 4.
-Given the description of $P$ above and assuming that $w_{1}$ is colored red, we have that
-  • the paths $v_{1},\dots,v_{i}$ (if they exist) and $x_{1},\dots,x_{i-1}$ are red-green colored paths,
-  • the paths $w_{i+1},\dots,w_{k-2}$ and $y_{i},\dots,y_{k-2}$ (if exists) are $\alpha$-green colored paths, where $\alpha$ is the color of $x_{k-2}$.
-Moreover, the endvertices of these paths $v_{1}$, $x_{1}$, $w_{k-2}$, $y_{k-2}$ have color green whereas the color of $v_{i},x_{i-1},w_{i+1},y_{i}$ depend on the parity of $i$ and $k.$
-Proof.
-The color of $x_{1}$ is green by maximality of $P$ and $x_{2}$ is red in a proper coloring, because $w_{2}$ is blue by our assumption.
-For each $3\leq i^{\prime}\leq i-1$, one can find the color of $x_{i^{\prime}}$ inductively: The color of $x_{i^{\prime}}$ is green if $i^{\prime}$ is odd, because $x_{i^{\prime}-1}$ has color red by the inductive hypothesis and $w_{i^{\prime}}$ has color red because of our assumption that $w_{1}$ is red.
-Note that if $x_{i^{\prime}}$ has color blue, we obtain a longer bicolored path by the union of $P$, $x_{i^{\prime}-1}$ and $x_{i^{\prime}}$, a contradiction.
-The color of $x_{i^{\prime}}$ is red if $i^{\prime}$ is even, because $x_{i^{\prime}-1}$ has color green by the inductive hypothesis and $w_{i^{\prime}}$ has color blue because of our assumption that $w_{1}$ is red.
-Also, if $C_{j-1}$ exists, that is $j\not=1$, we observe that $v_{1}$ is green by maximality of $P$ and $v_{1},v_{2},\dots,v_{i}$ is a red-green path by similar reasoning as above.
-Let $\alpha$ be the color of $x_{k-2},$ where $\alpha$ is either red or blue, depending on the modularity of $k.$ Let $\bar{\alpha}$ be the color on $P$ that is different than $\alpha,$ that is $\bar{\alpha}$ is blue if $\alpha$ is red and vice versa.
-The color of $w_{k-2}$ is green by maximality of $P$ and $w_{k-3}$ has color $\alpha$ in a proper coloring, since $x_{k-3}$ has color $\bar{\alpha}$.
-For each $4\leq i^{\prime}\leq k-i-1$, one can find the color of $w_{k-i^{\prime}}$ inductively: The color of $w_{k-i^{\prime}}$ is green if $i^{\prime}$ is even, since $w_{k-(i^{\prime}-1)}$ has color $\alpha$ by the inductive hypothesis and $x_{k-i^{\prime}}$ has color $\alpha$ as well.
-Note that if $w_{k-i^{\prime}}$ has color $\bar{\alpha}$, we obtain a longer bicolored path by the union of $P$, $w_{k-i^{\prime}}$ and $w_{k-(i^{\prime}-1)}$, a contradiction.
-Similarly, the color of $w_{k-i^{\prime}}$ is $\alpha$ if $i^{\prime}$ is odd, because $w_{k-(i^{\prime}-1)}$ is green by the inductive hypothesis and $x_{k-i^{\prime}}$ has color $\bar{\alpha}$.
-If $C_{j+2}$ exists, that is $j\leq k-4$, $y_{k-2}$ is green by maximality of $P$ and by similar reasoning we obtain that $y_{i+1},\dots,y_{k-2}$ is a $\alpha$-green path.
-As the color of $v_{i},x_{i-1},w_{i+1},y_{i}$ depend on the parity of $i$ and $k,$ the discussion on determining the color of these vertices is continued in the case analysis below.
+By symmetry, we assume w.l.o.g.
+that the jump of $P$ (across columns) is from left to right, specifically from $C_{j}$ to $C_{j+1}$ with $j\geq 2$ and at the $i$’th row with $2\leq i\leq k-2.$ Thus, $P$ is $w_{1},\dots,w_{i},x_{i},\dots,x_{k-2}.$ Note that $C_{j-1}$ exists since $j\geq 2$, and that $w_{i-1}$ and $x_{i-1}$ exist since $i\geq 2.$ By arguments as in Case 1, $v_{1}$ is (colored) green, $v_{2}$ is red, $v_{3}$ is green,…, $v_{i}$ is red or green (depending on the parity of $i$).
+The following can be proved similarly.
+Claim 4.
+$v_{1},\dots,v_{i}$ and $x_{1},\dots,x_{i-1}$ are green-red colored paths with $v_{1}$ and $x_{1}$ colored green.
+If $\alpha$ is the color of $x_{k-2}$, then $w_{k-2},\dots,w_{i+1}$ and $y_{k-2},\dots,y_{i}$ (if it exists) are green-$\alpha$ colored paths with $w_{k-2}$ and $y_{k-2}$ colored green.
 […]
-[scale=0.4]Case1_maximality.jpg
+[scale=0.4]Case1_maximality_new.jpg
 Figure 4.
-The neighboring columns of maximal red-blue colored $P$ (bold edges) and possible colorings: (a) $i\not\equiv k$ (mod 2) and $i$, even, (b) $i\equiv k$ (mod 2) and $i$, even, (c) $i\not\equiv k$ (mod 2) and i, odd, (d) $i\equiv k$ (mod 2) and $i$, odd.
-Case 2.1. $i\not\equiv k$ (mod 2) and $i$, even.
-If $j\leq k-4$, then $C_{j+2}$ exists and by Remark 4, there is a red-green $P_{2i+1}$ induced by $\{x_{1},\dots,x_{i},w_{1}$, $v_{1},\dots,v_{i}\}$ and a blue-green $P_{2(k-i-1)+1}$ induced by $\{w_{i},\dots,w_{k-2},x_{k-2},y_{i},\dots,y_{k-2}\}$, as shown in Figure 4 (a).
+The neighboring columns of maximal red-blue colored $P$ (bold edges) and possible colorings, respectively in (a)-(d): i) $k$, even: $i\equiv k$ or $i\not\equiv k$ (mod 2), ii) $k$, odd: $i\equiv k$ or $i\not\equiv k$ (mod 2).
+If $k$ is even and i=k2i=k-2, then $v_{1},w_{1},x_{1},\dots,x_{k-2}$ is a green-red colored path by Claim 4, contradicting the maximality of $P.$ If $k$ is even, i<k2i<k-2 and $i$ is even, then $w_{i+1}$ is colored red, which means $w_{1},\dots,w_{i+1},x_{i+1},\dots,x_{k-2}$ is also a red-blue colored path.
+Thus, if $k$ is even and i<k2i<k-2, we may assume that $i$ is odd.
+If $k$ is even, $i$ is odd, and i<k2i<k-2, then $w_{1},v_{1},\dots,v_{i},w_{i},\dots,w_{k-2}$ is a red-green colored path by Claim 4, contradicting the maximality of $P.$ Thus, we have the following.
+Claim 5.
+$k$ is odd.
+If $i$ is odd (i.e., $k$ and $i$ are both odd), then by Claim 4, $x_{i-1}$ is colored red and $w_{i+1}$ (if exists) is colored blue, and therefore the path $w_{1}\dots,w_{i-1},x_{i-1},x_{i},w_{i},w_{i+1},x_{i+1},\dots,x_{k-2}$ (or the part of it that exists) is a red-blue colored path, contradicting the maximality of $P.$ Thus, we have the following.
+Claim 6.
+$i$ is even (which also means that i<k2i<k-2).
+To complete the proof, it suffices to produce a contradiction assuming that $k$ is odd and $i$ is even.
+If $j\leq k-4$, then $C_{j+2}$ exists and by Claim 4, the path $x_{1},\dots,x_{i},w_{1}$, $v_{1},\dots,v_{i}$ is a red-green colored $P_{2i+1}$ and the path $w_{i},\dots,w_{k-2},x_{k-2},y_{i},\dots,y_{k-2}$ is a blue-green colored $P_{2(k-i-1)+1}$, as shown in Figure 4 (a).
 One of these bicolored paths has at least $k$ vertices, depending on the value of $i$.
 If j=k3j=k-3, then $C_{j+2}$ does not exist.
-In this case, it is not guaranteed to have a bicolored $P_{k}$ by the neighbors of $V(P)$ and itself.
+In this case, it is not guaranteed to have a bicolored $P_{k}$ only using vertices in $P$ and their neighbors.
 So, we use Lemma 2 to show the existence of a bicolored $P_{k}$.
-Let $C$ be the bicolored red-blue component containing $P.$ If $C$ contains any vertex from the top or bottom sides outside $P,$ then $C$ contains a red-blue $P_{k}.$ If $C$ contains any vertex from the left side outside $P,$ then the red-blue path in $C$ connecting $P$ to the left side together with $P$ yield a bicolored path with at least $k$ vertices.
+Let $C$ be the bicolored red-blue component containing $P.$ If $C$ contains any vertex from the top or bottom sides outside $P,$ then $C$ contains a red-blue $P_{k}.$ If $C$ contains any vertex from the left side, then the red-blue path in $C$ connecting $P$ to the left side together with $P$ yield a bicolored path with at least $k$ vertices.
 Thus, we assume that $C$ is not connected to any side of the grid (other than the right side) outside $P$.
 This implies that there exists a partial walk of $C$ from $x_{k-2}$ to $w_{1}$.
-However, Lemma 2 yields that $x_{k-2}$ and $w_{1}$ have the same color, a contradiction.
-Case 2.2. $i\equiv k$ (mod 2) and $i$, even.
-[…]
-Case 2.3. $i\not\equiv k$ (mod 2) and i, odd.
-If $j\not=1$, by Remark 4, the union of the two red-green paths $v_{1},v_{2},\dots,v_{i}$ and $w_{i},\dots,w_{k-2}$, together with $x_{k-2}$ induce a bicolored $P_{k}.$ If $j=1$, then $C_{j+2}$ exists.
-By Remark 4, $x_{1},\dots,x_{i-1}$ and $y_{i},\dots,y_{k-2}$ are red-green paths.
-If $y_{i-1}$ (black vertex in Figure 4 (c)) has color blue, then there is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},y_{i-1},y_{i},x_{i},$ $\dots,x_{k-2}\}.$ Otherwise, $y_{i-1}$ has color green and there is a red-green $P_{k}$ induced by $\{w_{1},x_{1},\dots x_{i-1},$ $y_{i-1},\dots,y_{k-2},x_{k-2}\}.$
-Case 2.4. $i\equiv k$ (mod 2) and $i$, odd.
-There is a red-blue $P_{k}$ induced by $\{w_{1},\dots,w_{i-1},x_{i-1},x_{i},w_{i},w_{i+1},x_{i+1},\dots,x_{k+2}\}.$
-As a special case, if $i=1$, then $P$ crosses between columns at the top row having vertex set $\{w_{1},x_{1},x_{2},\dots,x_{k-2}\}$.
-By Remark 4, $w_{2},\dots,w_{k-2}$ induce a red-green path and $k$ must be even, since $w_{k-2}$ is green by maximality of $P$.
+However, (4) of Lemma 2 yields that $x_{k-2}$ and $w_{1}$ have the same color, a contradiction.
+As a special case, if $i=1$, then $P$ is $w_{1},x_{1},x_{2},\dots,x_{k-2}$ and crosses between columns at the top row.
+By Claim 4, $w_{2},\dots,w_{k-2}$ is a red-green path and $k$ must be even, since $w_{k-2}$ is green by maximality of $P$.
 If $C_{j-1}$ exists, $v_{1}$ can be either green or blue in a proper coloring, inducing either a red-green $P_{k}$ together with $w_{1},\dots,w_{k-2},x_{k-2}$ or a red-blue $P_{k}$ together with $P$, respectively.
-In case, $j=1$, then $C_{j+2}$ exists and the path $y_{2},\dots,y_{k-2}$ is a red-green colored path by Remark 4.
-The red-green colored paths on $C_{j}$ and $C_{j+2}$ together $x_{k-2}$ yield a red-green $P_{k}.$ The same proof holds also for i=k2i=k-2 by symmetry.
+In case, $j=1$, $C_{j+2}$ exists and the path $y_{2},\dots,y_{k-2}$ is a red-green colored path by Claim 4.
+The red-green colored paths on $C_{j}$ and $C_{j+2}$ together $x_{k-2}$ yield a red-green $P_{k}.$
 Proof of Theorem 1: We know that $s_{k}(P_{m}\square P_{n})\leq 4$ as $s_{5}(P_{m}\square P_{n})=4$, for any $m,n\geq 3$, $k\geq 5$ ([18]).
@@
 Let $C$ be a partial, $c_{1}$-$c_{2}$ colored component containing vertices from the top side and possibly left side of the grid.
-Let $B_{C}=(v_{1},v_{2},\dots,v_{r})$ be a partial walk of $C$.
+Let $B_{C}=(v_{1},v_{2},\dots,v_{r})$ be a partial walk of $C$, where $v_{1}$ has color $c_{1}$.
 We need the following definitions associated with $C$:
-  • $D_{C}$: By Lemma 2, there is a $c_{1}$-$c_{3}$ colored connected subgraph, $c_{1}$ being the color of $v_{1}$, induced by $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}\not\in C$ is the green vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$ Let $D_{C}$ be the bicolored component containing this $c_{1}$-$c_{3}$ colored connected subgraph.
-  • $\mathcal{B}_{C}$: Let this be the collection of $C_{4}$’s induced by $\{u_{(i+1)/2},v_{i},v_{i+1},v_{i+2}\}$, for odd $i$, $1\leq i\leq r-2$. As $r\geq 3$, $\mathcal{B}_{C}\neq\emptyset$.
+  • $D_{C}$: By Lemma 2, there is a $c_{1}$-$c_{3}$ colored connected subgraph, induced by $v_{i}$ and $u_{(i+1)/2}$, for odd $i$, $1\leq i\leq r-2,$ where $u_{(i+1)/2}\not\in C$ is the $c_{3}$-colored vertex on the $C_{4}$ containing $\{v_{i},v_{i+1},v_{i+2}\}.$ Let $D_{C}$ be the bicolored component containing this $c_{1}$-$c_{3}$ colored connected subgraph.
+  • $\mathcal{W}_{C}$: Let this be the collection of $C_{4}$’s induced by $\{u_{(i+1)/2},v_{i},v_{i+1},v_{i+2}\}$, for odd $i$, $1\leq i\leq r-2$. As $r\geq 3$, $\mathcal{W}_{C}\neq\emptyset$.
   • $R_{C}$: This is the region of the grid enclosed by $B_{C}$ and the walk along the four sides of the grid in clockwise direction starting at $v_{1}$ ending at $v_{r}$.
@@
     Both $C$ and $D_{C}$ have some vertices from the top and left side of the grid.
-Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side of the grid, $v_{r}$, $v_{r}^{\prime}$ are on the left side of the grid.
+Thus, $v_{1}$, $v_{1}^{\prime}$ are on the top side of the grid, $v_{r}$, and $v_{r}^{\prime}$ are on the left side of the grid.
+[…]
+[width=0.5]main_iteration.jpg
+Figure 5.
+An example according to the scenario (2), the left and right grids showing the current $C$ before and after the iteration (and the flip), respectively.
+$\mathcal{R}_{C}$ is shown with the shaded region and $B_{C}$ is marked with bold edges.
 In these cases, the grid is not replaced with any of its automorphisms and the position of the vertices on the grid do not change at any iteration, except in (2).
@@
 Similarly, in all iterations in (1), $v_{r}^{\prime}$ is to the left of $v_{r}$ on the top side of the grid or $v_{r}=v_{r}^{\prime}$, and in (3), $v_{r}^{\prime}$ is below $v_{r}$ on the left side of the grid or $v_{r}=v_{r}^{\prime}$.
-Note that, after an iteration, the squares enclosed by the $C_{4}$’s in $\mathcal{B}_{C}$ are missing in $R_{C}$.
-As $\mathcal{B}_{C}\neq\emptyset$, $R_{C}$ becomes smaller after each iteration in cases (1)-(3).
+Note that, after an iteration, the squares enclosed by the $C_{4}$’s in $\mathcal{W}_{C}$ are missing in $R_{C}$.
+As $\mathcal{W}_{C}\neq\emptyset$, $R_{C}$ becomes smaller after each iteration in cases (1)-(3).
 Thus, as long as $C$ is not peripheral, $v_{1}=v_{1}^{\prime}$ or $v_{r}=v_{r}^{\prime}$ can happen at most once for each vertex, as there are exactly two $C_{4}$’s that contain $v_{1}$ (same for $v_{r}$) and no $C_{4}$ belongs to $R_{C}$ in more than one iteration.